OpenStudy (anonymous):

What is the radius of a circle represented by the given equation? x2 + y2 – 4x – 6y + 12 = 0?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

Huey Freeman?

OpenStudy (anonymous):

That Be me

OpenStudy (anonymous):

Cool :D

OpenStudy (anonymous):

Its either 5 or 12

OpenStudy (aum):

x^2 + y^2 – 4x – 6y + 12 = 0 Group the x-terms together and the y-terms together: (x^2 - 4x) + (y^2 - 6y) + 12 = 0 Complete the square of x^2 - 4x Complete the square of y^2 - 6y What do you get?

OpenStudy (anonymous):

12

OpenStudy (aum):

You need to complete the square first.

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Looks like you just want answers and I am sorry I can't do that. You should learn how to complete squares.

OpenStudy (anonymous):

I think I know what your talking about but I dont want to do anything wrong

OpenStudy (anonymous):

Im unsure

OpenStudy (anonymous):

(x-2)^2 +(y-3)^2 =12+ (-2)^2 +(-3)^2

OpenStudy (aum):

The 12 on the RHS should be -12.

OpenStudy (anonymous):

(x-2)^2 +(y-3)^2=

OpenStudy (aum):

(x-2)^2 +(y-3)^2 =-12+ (-2)^2 +(-3)^2 (x-2)^2 +(y-3)^2 =-12+ 4 + 9 = -12 + 13 = 1 (x-2)^2 +(y-3)^2 =1 = (1)^2 Compare it to: (x-h)^2 + (y-k)^2 = r^2 where (h, k) is the center of the circle and 'r' is the radius. They just want 'r' in this problem. r = ?

OpenStudy (anonymous):

1

OpenStudy (aum):

(x-2)^2 +(y-3)^2 = (1)^2 (x-h)^2 + (y-k)^2 = r^2 r = ?

OpenStudy (anonymous):

r=1

OpenStudy (anonymous):

?

OpenStudy (aum):

correct. The radius is 1.

OpenStudy (anonymous):

ok thx

OpenStudy (aum):

yw.